OpenStudy (anonymous):

Eliminate the parameter. x = t2 + 2, y = t2 - 4

OpenStudy (anonymous):

to eliminate find t^2 from one equation and substitute in other.

OpenStudy (anonymous):

May you guide me, I just got into this chapter and I really do not get it; please.

OpenStudy (anonymous):

y=t^2-4 take -4 to other side of equality

OpenStudy (anonymous):

CAN YOU WRITE IT ?

OpenStudy (anonymous):

y=t^2-4, y=t^2+4?

OpenStudy (anonymous):

t^2=y+4 now plug in other equation.

OpenStudy (anonymous):

t^2-2

OpenStudy (anonymous):

x=t^2+2 x=y+4+2 orx=?

OpenStudy (anonymous):

Shouldn't t^2 + 2 be t^2 - 2? y=2/4 + 2. I really don't know.

OpenStudy (anonymous):

y=2/4+2 is the answer I think it is.

OpenStudy (anonymous):

x =y+6 is the solution.

OpenStudy (anonymous):

With your answer I keep getting this question incorrect, so I'll just give you my options; y = x - 6, x ≥ 1 y = x + 6, x ≥ 1 y = x2 - 6, x ≥ 1 y = x2 + 6, x ≥ 1

OpenStudy (anonymous):

x=y+6 rearrange it y=x-6

OpenStudy (anonymous):

So, is this the correct one: y = x - 6, x ≥ 1?

OpenStudy (anonymous):

x should be >=2 but i don't see in your options. otherwise it is correct.

OpenStudy (anonymous):

It's correct. Thank You !